please help me to solve the problem
places A and B are 100km apart on a highway. one car starts from A and another starts from B . if the cars travel in the same directions with different speed , they meet in 5 hours . if they travel towards each other, they meet in 1 hour. what are the speeds of the two cars

Answer by stanbon(75887)

places A and B are 100km apart on a highway. one car starts from A and another starts from B . if the cars travel in the same directions with different speed , they meet in 5 hours . if they travel towards each other, they meet in 1 hour. what are the speeds of the two cars
---------------------------
Catch-up DATA:
A's DATA:
rate = a km/hr ; time = 5 hrs; distance = 5a km
B's DATA:
rate = b km/hr ; time = 5 hrs; distance = 5b km
---
Note If A catches up to B, A's distance is 100 km + 5b
So, 5a = 5b+100
------------------------
Meeting DATA:
A's DATA:
rate = a km/hr ; time = 1 hr; distance = a km
B's DATA:
rate = b km/hr ; time = 1 hr; distance = b km
---
Equation:
a+b = 100 km
-----------------------
So you have two equations:
5a = 5b+100
a+b = 100
----
Modifying those you get:
a-b = 20
a+b = 100
----
Add and solve for "a:
2a = 120
a = 60 km/hr (A's rate)
---
Solve for "b":
a + b = 100
60 +b = 100
b = 40 mph (B's rate)
==========================
Cheers,
Stan H.